### v3

# Quantum Speedups in Regret Analysis of Infinite Horizon Average-Reward Markov Decision Processes

Abstract
[…]
1 Introduction
Quantum Machine Learning (QML) has garnered remarkable interest in contemporary research, predominantly attributed to the pronounced speedups achievable with quantum computers as opposed to their classical analogs (?, ?). The fundamental edge of quantum computing arises from the unique nature of its fundamental computing element, termed a qubit, which can exist simultaneously in states of 0 and 1, unlike classical bits that are restricted to either 0 or 1. This inherent distinction underpins the exponential advancements that quantum computers bring to specific computational tasks, surpassing the capabilities of classical computers.
In the realm of Reinforcement Learning (RL), an agent embarks on the task of finding an efficient policy for Markov Decision Process (MDP) environment through repetitive interactions (?). RL has found notable success in diverse applications, including but not limited to autonomous driving, ridesharing platforms, online ad recommendation systems, and proficient gameplay agents (?, ?, ?, ?). Most of these setups require decision making over infinite horizon, with an objective of average rewards. This setup has been studied in classical reinforcement learning (?, ?, ?), where $\tilde{\mathcal{O}}(\sqrt{T})$ regret across $T$ rounds has been found meeting the theoretical lower bound (?). This paper embarks on an inquiry into the potential utilization of quantum statistical estimation techniques to augment the theoretical convergence speeds of tabular RL algorithms within the context of infinite horizon learning settings.
A diverse array of quantum statistical estimation primitives has emerged, showcasing significant enhancements in convergence speeds and gaining traction within Quantum Machine Learning (QML) frameworks (?, ?, ?). Notably, (?) introduces a quantum mean estimation algorithm that yields a quadratic acceleration in sample complexity when contrasted with classical mean estimation techniques. In this work, we emphasize the adoption of this particular mean estimation method as a pivotal component of our methodology. It strategically refines the signals garnered by the RL agent during its interaction with the enigmatic quantum-classical hybrid MDP environment.
It is pertinent to highlight a fundamental element in the analysis of conventional Reinforcement Learning (RL), which involves the utilization of martingale convergence theorems. These theorems play a crucial role in delineating the inherent stochastic process governing the evolution of states within the Markov Decision Process (MDP). Conversely, this essential aspect lacks parallelism within the quantum setting, where comparable martingale convergence results remain absent. To address this disparity, we introduce an innovative approach to regret bound analysis for quantum RL. Remarkably, our methodology circumvents the reliance on martingale concentration bounds, a staple in classical RL analysis, thus navigating uncharted territories in the quantum realm.
Moreover, it is important to highlight that mean estimation in quantum mechanics results in state collapse, which makes it challenging to perform estimation of state transition probabilities over multiple epochs. In this study, we present a novel approach to estimating state transition probabilities that explicitly considers the quantum state collapsing upon measurement.
To this end, the following are the major contributions of this work:
  1. 1.
    Our paper introduces a pioneering contribution in the form of Quantum-UCRL (Q-UCRL), an innovative model-based, infinite horizon, optimism-driven quantum Reinforcement Learning (QRL) algorithm. Drawing inspiration from classical predecessors like the UCRL and UC-CURL algorithms (?, ?, ?), Q-UCRL is designed to seamlessly weave together an agent’s optimistic policy acquisition and an adept quantum mean estimator.
  2. 2.
    Employing meticulous theoretical analysis, we show that the Q-UCRL algorithm achieves an exponential improvement in the regret analysis. Specifically, it attains a regret bound of $\tilde{\mathcal{O}}(1)$, breaking through the theoretical classical lower bound of $\Omega(\sqrt{T})$ across $T$ online rounds. The analysis is based on the novel quantum Bellman error based analysis (introduced in classical RL in (?) for optimism based algorithm), where the difference between the performance of a policy on two different MDPs is bounded by long-term averaged Bellman error and the quantum mean estimation is used for improved Bellman error. Further, the analysis avoids dependence on martingale concentration bounds and incorporates the phenomenon of state collapse following measurements, a crucial aspect in quantum computing analysis.
To the best of our knowledge, these are the first results for quantum speedups for infinite horizon MDPs with average reward objective.
2 Related Work
Infinite Horizon Reinforcement Learning: Regret analysis of infinite horizon RL in the classical setting has been widely studied with average reward objective. A prominent principle that underpins algorithms tailored for this scenario is the concept of “optimism in the face of uncertainty” (OFU). In this approach, the RL agent nurtures optimistic estimations of value functions and, during online iterations, selects policies aligned with the highest value estimates (?, ?, ?). Additionally, it’s noteworthy to acknowledge that several methodologies are rooted in the realm of posterior sampling, where the RL agent samples an MDP from a Bayesian Distribution and subsequently enacts the optimal policy (?, ?, ?, ?). In our study, we embrace the OFU-based algorithmic framework introduced in (?), and we extend its scope to an augmented landscape where the RL agent gains access to supplementary quantum information. Furthermore, we render a mathematical characterization of regret, revealing a $\tilde{\mathcal{O}}(1)$ bound, which in turn underscores the merits of astutely processing the quantum signals within our framework.
Quantum Mean Estimation: The realm of mean estimation revolves around the identification of the average value of samples stemming from an unspecified distribution. Of paramount importance is the revelation that quantum mean estimators yield a quadratic enhancement when juxtaposed against their classical counterparts (?, ?). The key reason for this improvement is based on the quantum amplitude amplification, which allows for suppressing certain quantum states w.r.t. the states that are desired to be extracted (?). In Appendix B, we present a discussion around Quantum Amplitude Estimation and its applications in QML.
In the following, we introduce the definition of key elements and results pertaining to quantum mean estimation that are critical to our setup and analysis. First, we present the definition of a classical random variable and the quantum sampling oracle for performing quantum experiments.
Definition 1 (Random Variable, Definition 2.2 of (?)).
A finite random variable can be represented as $X:\Omega\to E$ for some probability space $(\Omega,\mathbb{P})$, where $\Omega$ is a finite sample set, $\mathbb{P}:\Omega\to[0,1]$ is a probability mass function and $E\subset\mathbb{R}$ is the support of $X$. $(\Omega,\mathbb{P})$ is frequently omitted when referring to the random variable $X$.
To perform quantum mean estimation, we provide the definition of quantum experiment. This is analogous to the classical random experiments
Definition 2 (Quantum Experiment).
Consider a random variable $X$ on a probability space $(\Omega,2^{\Omega},\mathbb{P})$. Let $\mathcal{H}_{\Omega}$ be a Hilbert space with basis states $\{\lvert\omega\rangle\}_{\omega\in\Omega}$ and fix a unitary $\mathcal{U}_{\mathbb{P}}$ acting on $\mathcal{H}_{\Omega}$ such that
$\mathcal{U}_{\mathbb{P}}:\lvert 0\rangle\mapsto\sum_{\omega\in\Omega}\sqrt{\mathbb{P}(\omega)}\lvert\omega\rangle$
assuming $0\in\Omega$. We define a quantum experiment as the process of applying the unitary $\mathcal{U}_{\mathbb{P}}$ or its inverse $\mathcal{U}_{\mathbb{P}}^{-1}$ on any state in $\mathcal{H}_{\Omega}$.
The unitiary $\mathcal{U}_{\mathbb{P}}$ provides the ability to query samples of the random variable in superpositioned states, which is the essence for speedups in quantum algorithms. To perform quantum mean estimation for values of random variables (?, ?, ?), an additional quantum evaluation oracle would be needed.
Definition 3 (Quantum Evaluation Oracle).
Consider a finite random variable $X:\Omega\to E$ on a probability space $(\Omega,2^{\Omega},\mathbb{P})$. Let $\mathcal{H}_{\Omega}$ and $\mathcal{H}_{E}$ be two Hilbert spaces with basis states $\{|\omega\rangle\}_{\omega\in\Omega}$ and $\{|x\rangle\}_{x\in E}$ respectively. We say that a unitary $\mathcal{U}_{X}$ acting on $\mathcal{H}_{\Omega}\otimes\mathcal{H}_{E}$ is a quantum evaluation oracle for $X$ if
$\mathcal{U}_{X}:|\omega\rangle|0\rangle\mapsto|\omega\rangle|X(\omega)\rangle$
for all $\omega\in\Omega$, assuming $0\in E$.
Next, we present a key quantum mean estimation result that will be carefully employed in our algorithmic framework to utilize the quantum superpositioned states collected by the RL agent. One of the crucial aspects of Lemma 1 is that quantum mean estimation converges at the rate $\mathcal{O}(\frac{1}{n})$ as opposed to classical benchmark convergence rate $\mathcal{O}(\frac{1}{\sqrt{n}})$ for $n$ number of samples, therefore estimation efficiency quadratically.
Lemma 1 (Quantum multivariate bounded estimator, Theorem 3.3 of (?)).
Let $X$ be a d-dimensional bounded random variable such that $||X||_{2}\leq 1$. Given three reals $L_{2}\in(0,1]$, $\delta\in(0,1)$ and $n\geq 1$ such that $\mathbb{E}[||X||_{2}]\leq L_{2}$, the multivariate bounded estimator $\texttt{QBounded}_{d}(X,L_{2},n,\delta)$ obtained by Algorithm 1 outputs a mean estimate $\hat{\mu}$ of $\mu=\mathbb{E}[X]$ such that,
${\mathbb{P}\left[||\hat{\mu}-\mu||_{\infty}\leq\frac{\sqrt{L_{2}}\log(d/\delta)}{n}\right]\geq 1-\delta.}$ (1)
The algorithm performs $\mathcal{O}\Big{(}n\log^{1/2}(n\sqrt{d})\Big{)}$ quantum experiments.
Algorithm 1 $\texttt{QBounded}_{d}(X,L_{2},n,\delta)$ (Algorithm 1 in (?))
1:if $n\leq\frac{\log(d/\delta)}{\sqrt{L_{2}}}$ then
2:     Output $\hat{\mu}=0$.
3:end if
4:Set $\alpha=\frac{1}{\sqrt{\log(400\pi n\sqrt{d})}}$ and $m=2^{\left\lceil\log\left(\frac{8\pi n}{\alpha\sqrt{L_{2}\log(d/\delta)}}\right)\right\rceil}$.
5:Set $G=\left\{\frac{j}{m}-\frac{1}{2}+\frac{1}{2m}:j\in\{0,\dots,m-1\}\right\}^{d}\subseteq\left(-\frac{1}{2},\frac{1}{2}\right)^{d}$
6:for $k=1,\dots,\lceil 18\log(d/\delta)\rceil$ do
7:     Compute the uniform superposition $|G\rangle:=\frac{1}{\sqrt{md/2}}\sum_{u\in G}|u\rangle$ over $G$.
8:     Compute the state $|\psi\rangle:=\widetilde{P}_{X,L_{2},m,\alpha,\epsilon}|G\rangle$ in $H_{G}\otimes H_{aux}$, where $\widetilde{P}_{X,L_{2},m,\alpha,\epsilon}$ is the directional mean oracle constructed by the quantum evaluation oracle in Proposition 3.2 (?) with $\epsilon=1/25$.
9:     Compute the state $|\phi\rangle:=(QFT_{G}^{-1}\otimes I_{aux})|\psi\rangle$ where the unitary $QFT_{G}:|u\rangle\mapsto\frac{1}{\sqrt{md/2}}\sum_{v\in G}e^{2\pi i\langle u,v\rangle}|v\rangle$ is the quantum Fourier transform over $G$.
10:     Measure the $H_{G}$ register of $|\phi\rangle$ in the computational basis and let $v^{(k)}\in G$ denote the obtained result. Set $\mu^{(k)}=\frac{2\pi v^{(k)}}{\alpha}$.
11:end for
12:Output the coordinate-wise median $\hat{\mu}=\text{median}(\mu^{(1)},\dots,\mu^{(\lceil 18\log(d/\delta)\rceil)})$.
Quantum Reinforcement Learning: Within the realm of QRL, a prominent strand of previous research showcases exponential enhancements in regret through amplitude amplification-based methodologies applied to the Quantum Multi-Armed Bandits (Q-MAB) problem (?, ?). Nevertheless, the theoretical underpinnings of the aforementioned Q-MAB approaches do not seamlessly extend to the QRL context since there is no state evolution in bandits.
A recent surge of research interest has been directed towards Quantum Reinforcement Learning (QRL) (?, ?, ?). It is noteworthy to emphasize that the aforementioned studies do not provide an exhaustive mathematical characterization of regret. (?) delves into QRL within the infinite horizon Markov Decision Process (MDP) learning framework, yet their focus rests on a discounted reward setup, where advancements in sample complexity are demonstrated through the lens of generative model assumptions. Further, finite horizon episodic MDP setup has been considered in (?). Ours is the first paper to study infinite horizon MDP with average reward objective.
3 Problem Formulation
We consider the problem of Reinforcement Learning (RL) in an infinite horizon Markov Decision Process characterized by $\mathcal{M}\triangleq(\mathcal{S},\mathcal{A},P,r,D)$, wherein $\mathcal{S}$ and $\mathcal{A}$ represent finite collections of states and actions respectively with $|\mathcal{S}|=S$ and $|\mathcal{A}|=A$, pertaining to RL agent’s interaction with the unknown MDP environment. $P(s^{\prime}|s,a)\in[0,1]$ denotes the transition probability for next state $s^{\prime}\in\mathcal{S}$ for a given pair of previous state and RL agent’s action, i.e., $(s,a)\in\mathcal{S}\times\mathcal{A}$. Further, $r:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ represents the reward collected by the RL agent for state-action pair $(s,a)$. $D$ is the diameter of the MDP $\mathcal{M}$. In the following, we first present the additional quantum transition oracle that the agent could access during its interaction with the unknown MDP environment at every RL round. Here, we would like to emphasize that unknown MDP environment implies that matrix $P$ which encapsulates the transition dynamics of the underlying RL environment is not known beforehand.
Quantum Computing Basics: In this section, we give a brief introduction to the concepts that are most related to our work based on (?) and (?). Consider an $m$ dimensional Hilbert space $\mathbb{C}^{m}$, a quantum state $|x\rangle=(x_{1},...,x_{m})^{T}$ can be seen as a vector inside the Hilbert space with $\sum_{i}|x_{i}|^{2}=1$. Furthermore, for a finite set with $m$ elements $\xi=\{\xi_{1},...,\xi_{m}\}$, we can assign each $\xi_{i}\in\xi$ with a member of an orthonormal basis of the Hilbert space $\mathbb{C}^{m}$ by
$\xi_{i}\mapsto|\xi_{i}\rangle\equiv e_{i}$ (2)
where $e_{i}$ is the $i$th unit vector for $\mathbb{C}^{m}$. Using the above notation, we can express any arbitrary quantum state $|x\rangle=(x_{1},...,x_{m})^{T}$ by elements in $\xi$ as:
$|x\rangle=\sum_{n=1}^{m}x_{n}|\xi_{n}\rangle$ (3)
where $|x\rangle$ is the quantum superposition of the basis $|\xi_{1}\rangle,...,|\xi_{m}\rangle$ and we denote $x_{n}$ as the amplitude of $|\xi_{n}\rangle$. To obtain a classical information from the quantum state, we perform a measurement and the quantum state would collapse to any basis $|\xi_{i}\rangle$ with probability $|x_{i}|^{2}$. In quantum computing, the quantum states are represented by input or output registers made of qubits that could be in superpositions.
Quantum transition oracle: We utilize the concepts and notations of quantum computing in (?, ?, ?, ?) to construct the quantum sampling oracle of RL environments and capture agent’s interaction with the unknown MDP environment. We now formulate the equivalent quantum-accessible RL environments for our classical MDP $\mathcal{M}$. For an agent at step $t$ in state $s_{t}$ and with action $a_{t}$, we construct the quantum sampling oracles for variables of the next state $P(\cdot|s_{t},a_{t})$. Specifically, suppose there are two Hilbert spaces $\mathcal{\bar{S}}=\mathbb{C}^{|\mathcal{S}|}$ and $\mathcal{\bar{A}}=\mathbb{C}^{|\mathcal{A}|}$ containing the superpositions of the classical states and actions. We represent the computational bases for $\mathcal{\bar{S}}$ and $\mathcal{\bar{A}}$ as $\{|s\rangle\}_{s\in\mathcal{S}}$ and $\{|a\rangle\}_{a\in\mathcal{A}}$. We assume that we can implement the quantum sampling oracle $\mathcal{U}_{P}$ of the MDP’s transitions as follows:
  • The quantum evaluation oracle for the transition probability (quantum transition oracle) $\mathcal{U}_{P}$ which at step $t$, returns the superposition over $s^{\prime}\in\mathcal{S}$ according to $P(s^{\prime}|s_{t},a_{t})$, the probability distribution of the next state given the current state $|s_{t}\rangle$ and action $|a_{t}\rangle$ is defined as:
    $\mathcal{U}_{P}:|s_{t}\rangle\otimes|a_{t}\rangle\otimes|0\rangle\rightarrow|\bm{\psi}\rangle^{t}$
    Where $|\bm{\psi}\rangle^{t}=|s_{t}\rangle\otimes|a_{t}\rangle\otimes\sum_{s^{\prime}\in\mathcal{S}}\sqrt{P(s^{\prime}|s_{t},a_{t})}|s^{\prime}\rangle$.
Figure 1: Agent’s interaction at round $t$ with the MDP Environment and accessible quantum transition oracle
As described in Fig 1, at step $t$, the agent plays an action $a_{t}$ based on current state $s_{t}$ according to its policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$. Consequently, the unknown MDP Q-environment shares the next state information and reward i.e., $\{s_{t+1},r_{t}\}$. Additionally, the quantum transition oracle first encodes $s_{t}$ and $a_{t}$ into the corresponding basis of the quantum Hilbert spaces, producing basis $|s_{t}\rangle$ and $|a_{t}\rangle$, then it encodes the superpositioned quantum state $|\bm{\psi}\rangle^{t}$ from the quantum transition oracle $\mathcal{U}_{P}$. $|\bm{\psi}\rangle^{t}$ would be used for the Quantum Mean Estimator QBounded outlined in Lemma 1 which in turn improves transition probability estimation and leads to exponentially better regret. We note that a single copy of $|\bm{\psi}\rangle^{t}$ is obtained at each $(s_{t},a_{t})$. Given that a measurement collapses the state, we can only perform one measurement on it. Thus, the quantum mean estimator will only perform measurements at the end of epochs, as will be described in the algorithm.
RL regret formulation: Consider the RL agent following some policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ on the unknown MDP Q-environment $\mathcal{M}$. Then, as a result of playing policy $\pi$, the agent’s long term average reward is defined next according to (?).
$\Gamma_{\pi}^{P}=\lim_{T\rightarrow\infty}\mathbb{E}_{\pi,P}\big{[}\frac{1}{T}\sum_{t=0}^{T-1}~{}r(s_{t},a_{t})\big{]},$ (4)
where $\mathbb{E}_{\pi,P}[\cdot]$ is the expectation taken over the trajectory $\{(s_{t},a_{t})\}_{t\in[0,T-1]}$ by playing policy $\pi$ on $\mathcal{M}$ at every time step. In this context, we state the definitions of value function for MDP $\mathcal{M}$ next:
$V_{\pi}^{{P}}(s;\gamma)$ $=\mathbb{E}_{a\sim\pi}\Big{[}Q_{\pi}^{{P}}(s,a;\gamma)\Big{]},$ (5)
$=\mathbb{E}_{a\sim\pi}\Big{[}r(s,a)+\gamma\sum_{s^{\prime}\in\mathcal{S}}P(s^{\prime}|s,a)V_{\pi}^{{P}}(s^{\prime};\gamma)\Big{]}$ (6)
Then, according to (?), $\Gamma_{\pi}^{P}$ can be represented as the following:
$\Gamma_{\pi}^{P}$ $=\lim_{\gamma\rightarrow 1}(1-\gamma)V_{\pi}^{{P}}(s;\gamma),~{}\forall s\in\mathcal{S},$ (7)
$=\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\rho_{\pi}^{P}(s,a)\overline{r}(s,a),$ (8)
wherein, $V_{\pi}^{P}(\cdot;\cdot)$ is the discounted cumulative average reward by following policy $\pi$, $\rho_{\pi}^{P}$ is the steady-state occupancy measure (i.e., $\rho_{\pi}^{P}(s,a)\in[0,1],~{}\sum_{s\in\mathcal{S}}\sum_{a\in\mathcal{A}}\rho_{\pi}^{P}(s,a)=1$), $\bar{r}(s,a)$ is the steady-state long-term average reward of pair $(s,a)$. In the following, we introduce key assumptions crucial from the perspective of our algorithmic framework and its analysis. In this context, we introduce the notations $\{P_{\pi,s}^{t},T_{\pi,~{}s\rightarrow s^{\prime}}\}$ to denote the $t$ time step probability distribution obtained by playing policy $\pi$ on the unknown MDP $\mathcal{M}$ starting from some arbitrary state $s\in\mathcal{S}$; and, the actual time steps to reach $s^{\prime}$ from $s$ by playing policy $\pi$ respectively. With these notations, we introduce our first assumption pertaining to ergodicity of MDP $\mathcal{M}$.
Assumption 1 (Finite MDP mixing time).
We assume that unknown MDP environment $\mathcal{M}$ has finite mixing time, which mathematically implies:
$T_{\texttt{mix}}\triangleq\max_{\pi,(s,s^{\prime})\in\mathcal{S}\times\mathcal{S}}~{}\mathbb{E}[T_{\pi,~{}s\rightarrow s^{\prime}}]<\infty,$ (9)
where $T_{\pi,~{}s\rightarrow s^{\prime}}$ implies the number of time steps to reach a state $s^{\prime}$ from an initial state $s$ by playing some policy $\pi$.
Assumption 2.
The reward function $r(\cdot,\cdot)$ is known to the RL agent.
We emphasize that Assumption 2 is a commonly used assumption in RL algorithm development owing to the fact that in most setups rewards are constructed according to the underlying problem and is known beforehand (?, ?). Next, we define the cumulative regret accumulated by the agent in expectation across $T$ time steps as follows:
$\mathcal{R}_{[0,T-1]}\triangleq T.\Gamma_{\pi^{*}}^{P}-\mathbb{E}\Big{[}\sum_{t=0}^{T-1}r(s_{t},a_{t})\Big{]},$ (10)
where the agent generates the trajectory $\{(s_{t},a_{t}\}_{t\in[0,T-1]}$ by playing policy $\pi$ on $\mathcal{M}$, and $\pi^{*}$ is the optimal policy that maximizes the long term average reward defined in Eq. (4). Finally, the RL agent’s goal is to determine a policy $\pi$ which minimizes $\mathcal{R}_{[0,T-1]}$.
4 Algorithmic Framework
In this section, we first delineate the key intuition behind our methodology followed by formal description of Q-UCRL algorithm. If the agent was aware of the true transition probability distribution $P$, then it would simply solve the following optimization problem for the optimal policy.
$\max_{\{\rho(s,a)\}_{(s,a)\in\mathcal{S}\times\mathcal{A}}}~{}\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}r(s,a)\rho(s,a)$ (11)
With the following constraints,
$\sum_{s,a}\rho(s,a)=1,\rho(s,a)\geq 0\quad\forall(s,a)\in\mathcal{S}\times\mathcal{A}$ (12)
$\sum_{a\in\mathcal{A}}\rho(s^{\prime},a)=\sum_{s,a}P(s^{\prime}|s,a)\rho(s,a),\forall s^{\prime}\in\mathcal{S}$ (13)
Note that the objective in Eq. (11) resembles the definition of steady-state average reward formulation in Eq. (4). Furthermore, Eq. (12), (13) preserve the properties of the transition model. Consequently, once it obtains $\{\rho^{*}(s,a)\}_{(s,a)\in\mathcal{S}\times\mathcal{A}}$ by solving Eq. (11), it can compute the optimal policy as:
$\pi^{*}(a|s)=\frac{\rho^{*}(s,a)}{\sum_{\tilde{a}\in\mathcal{A}}\rho^{*}(s,\tilde{a})}\quad\forall s,a$ (14)
However, in the absence of knowledge pertaining to true $P$, we propose to solve an optimistic version of Eq. (11) that will update RL agent’s policy in epochs.
Remark 1.
Algorithm 2 runs in epochs wherein the agent’s action policy $\pi$ is updated only at the beginning of a new epoch, and the policy is held constant throughout the epoch.
Specifically, at the start of an epoch $e$, the following optimization problem will be solved:
$\max_{\{\rho(s,a),P_{e}(\cdot|s,a)\}_{(s,a)\in\mathcal{S}\times\mathcal{A}}}~{}\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}r(s,a)\rho(s,a),$ (15)
With the following constraints,
$\sum_{s,a}\rho(s,a)=1,\rho(s,a)\geq 0\quad\forall(s,a)\in\mathcal{S}\times\mathcal{A}$ (16)
$\sum_{s^{\prime}\in\mathcal{S}}P_{e}(s^{\prime}|s,a)=1,P_{e}(\cdot|s,a)\geq 0~{}\forall(s,a)\in\mathcal{S}\times\mathcal{A}$ (17)
$\sum_{a\in\mathcal{A}}\rho(s^{\prime},a)=\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}{P}_{e}(s^{\prime}|s,a)\rho(s,a),\forall s^{\prime}\in\mathcal{S}$ (18)
$\|{P}_{e}(\cdot|s,a)-\hat{P}_{e}(\cdot|s,a)\|_{1}\leq\frac{7S\log(S^{2}At_{e})}{\max\{1,N_{e}(s,a)\}}$ (19)
Observe that Eq. (15) increases the search space for policy in a carefully computed neighborhood of certain transition probability estimates $\hat{P}$ via Eq. (19) at the start of epoch $e$. Also in Eq. (19), recall that $S=|\mathcal{S}|$ is the number of states, $A=|\mathcal{A}|$ is the number of actions, $N_{e}(s,a)$ is the number of visitations of state-action pair $(s,a)$ upto the end of epoch $e$.
Note that, solving Eq. (15) is equivalent to obtaining an optimistic policy as it extends the search space of original problem Eq. (11). Furthermore, the computation of $\hat{P}_{e}$ is key to our framework design. Since performing quantum mean estimation to quantum samples would require measuring and results in quantum samples collapsing, each quantum sample can only be used once. We emphasize that the quantum mean estimator will only run at the end of the epoch, performing measurements to update the transition probability for the next epoch.
Details on Quantum Mean Estimator: At the end of the epoch, we use all $(s_{i},a_{i})$ observed in the samples, and perform quantum mean estimation on all the samples $|\bm{\psi}\rangle^{t}$ obtained for that $(s_{i},a_{i})$. More precisely, $\tilde{P}_{e}(\cdot|s_{i},a_{i})$ is the quantum mean estimation for the random variable $P_{e}(\cdot|s_{i},a_{i})$ performed by QBounded (Algorithm 1 in (?)) using the quantum samples $\{|\bm{\psi}\rangle\}^{e}_{i}$ observed in the previous epoch with the total number being $\nu_{e-1}(s_{i},a_{i})$:
${\{|\bm{\psi}\rangle\}^{e}_{i}=\{|\bm{\psi}\rangle^{t}\,|\;t_{e-1}^{start}\leq t\leq t_{e}^{start}-1,(s_{t},a_{t})=(s_{i},a_{i})\}}$ (20)
where $t_{e}^{start}$ is the starting step of epoch $e$. Using these samples, the transition probability is obtained as
${\tilde{P}_{e}(\cdot|s_{i},a_{i})=\texttt{QBounded}_{S}(\{P_{e}(\cdot|s_{i},a_{i}),L_{2},n_{e}^{s_{i},a_{i}},\delta_{e})}$ (21)
For a specific state-action pair $(s_{i},a_{i})$, we construct an estimator $\hat{P}_{e}(\cdot|s_{i},a_{i})$ used in epoch $e$ that consists of a weighted average of the estimator $\hat{P}_{e-1}(\cdot|s_{i},a_{i})$ and the estimator obtained using the quantum samples in the latest completed epoch $\tilde{P}_{e}(\cdot|s_{i},a_{i})$ as follows:
$\hat{P}_{e}(\cdot|s_{i},a_{i})=\begin{cases}0&\text{if $e=1$}\\ \tilde{P}_{e}(\cdot|s_{i},a_{i})&\text{if $e=2$}\\ \hat{w}\hat{P}_{e-1}(\cdot|s_{i},a_{i})+\tilde{w}\tilde{P}_{e}(\cdot|s_{i},a_{i})&\text{if $e\geq 3$}\\ \end{cases}$ (22)
Where
$\hat{w}=\frac{N_{e-1}(s_{i},a_{i})}{N_{e-1}(s_{i},a_{i})+\nu_{e-1}(s_{i},a_{i})}$ (23)
$\tilde{w}=\frac{\nu_{e-1}(s_{i},a_{i})}{N_{e-1}(s_{i},a_{i})+\nu_{e-1}(s_{i},a_{i})}$ (24)
In particular, given that $\nu_{e-1}(s_{i},a_{i})$ is the maximum number of quantum experiments we can perform in Eq. (21), we set
$n_{e}^{s_{i},a_{i}}=\left\lfloor\frac{\nu_{e-1}(s_{i},a_{i})}{c\log^{1/2}(T\sqrt{S})}\right\rfloor$ (25)
for some $c\in\mathbb{R}$ such that the term $cn\log^{1/2}(n\sqrt{S})$ is greater than the number of quantum experiments needed for performing Eq. (21), which is $\mathcal{O}\Big{(}n\log^{1/2}(n\sqrt{S})\Big{)}$. In addition, since $\mathbb{E}[||P_{e}(\cdot|s_{i},a_{i})||_{2}]\leq 1$, we set $L_{2}=1.$ Different from the works in classical approaches based on (?), our algorithm utilizes each quantum sample only once for estimating all $\hat{P}_{e}$ throughout the entire process. This strategy fits with our quantum framework since quantum samples would collapse once being measured. In Appendix A, we show that Eq. (15) is a convex optimization problem, therefore standard optimization solvers can be employed to solve it. Finally, the agent’s policy at the start of epoch $e$ is given by:
$\pi_{e}(a|s)=\frac{\rho_{e}(s,a)}{\sum_{\tilde{a}\in\mathcal{A}}\rho_{e}(s,\tilde{a})}$ (26)
Next, we formally present our method Q-UCRL in Algorithm 2.
Algorithm 2 Q-UCRL
1:Inputs: $\mathcal{S}$, $\mathcal{A}$, $r(\cdot,\cdot)$.
2:Set $t\leftarrow 1,~{}e\leftarrow 1,~{}\delta_{e}\leftarrow 0,t_{e}\leftarrow 0,t_{e}^{start}\leftarrow 1$
3:Set $\mu(s,a,s^{\prime})\leftarrow 0~{}\forall(s,a,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}$.
4:Set $\hat{P}_{e}(\cdot|s,a)\leftarrow 0~{}\forall(s,a)\in\mathcal{S}\times\mathcal{A}$.
5:for $(s,a)\in\mathcal{S}\times\mathcal{A}$ do
6:     Set $\nu_{e}(s,a)\leftarrow 0,~{}N_{e}(s,a)\leftarrow 0$.
7:end for
8:Obtain $\pi_{e}$ by solving Eq. (15) and Eq. (26).
9:for $t=1,2,\dots$ do
10:     Observe state $s_{t}$ and play action $a_{t}\sim\pi_{e}(\cdot|s_{t})$.
11:     Observe $s_{t+1},r(s_{t},a_{t})$ and quantum samples $|\bm{\psi}\rangle^{t}$.
12:     Update $\nu_{e}(s_{t},a_{t})\leftarrow\nu_{e}(s_{t},a_{t})+1$.
13:     Update $\mu(s_{t},a_{t},s_{t+1})\leftarrow\nu_{e}(s_{t},a_{t},s_{t+1})+1$.
14:     Set $t_{e}\leftarrow t_{e}+1$.
15:     if $\nu_{e}(s_{t},a_{t})=\max\{1,N_{e}(s_{t},a_{t})\}$ then
16:         for $(s,a)\in\mathcal{S}\times\mathcal{A}$ do
17:              $N_{e+1}(s,a)\leftarrow N_{e}(s,a)+\nu_{e}(s,a)$.
18:         end for
19:         Set $e\leftarrow e+1,~{}\nu_{e}(s,a)\leftarrow 0,~{}\delta_{e}\leftarrow\frac{1}{S^{2}At_{e}^{7}}$.
20:         Set $t_{e}\leftarrow 0,\;t_{e}^{start}\leftarrow t+1$.
21:         Obtain $\hat{P}_{e}(\cdot|s,a)$ by Eq. (22) $\forall(s,a)\in\mathcal{S}\times\mathcal{A}$
22:         Obtain $\pi_{e}$ by solving Eq. (15) and Eq. (26).
23:     end if
24:end for
Algorithm 2 proceeds in epochs, and each epoch $e$ contains multiple rounds of RL agent’s interaction with the environment. At each $t$, by playing action $a_{t}$ for insantaneous state $s_{t}$ through current epoch policy $\pi_{e}$, the agent collects next state configuration consisting of classical signals $s_{t+1},r(s_{t},a_{t})$ and a quantum sample $|\bm{\psi}\rangle^{(t)}$. Additionally, the agent keeps track of its visitations to the different states via the variables $\{\nu_{s,a},\mu_{s,a,s^{\prime}}\}$. Note that a new epoch is triggered at the “if” block of Line 15 in algorithm 2, which in turn is aligned to the widely used concept of doubling trick used in model-based RL algorithm design. Finally, the empirical visitation counts and probabilities up to epoch $e$, i.e., $\{N_{e+1}(s,a),\hat{P}_{e+1}(s^{\prime}|s,a)\}$ attributes are updated (line 15-21), and the policy is updated via re-solving from Eq. (15) - (19) and using Eq. (26).
5 Theoretical Results
In this section, we characterize the cumulative regret for $T$ rounds by running Q-UCRL in a unknown MDP hybrid Q-Environment. In the following, we first present a crucial result bounding the error accumulated till $e$ epochs between the true transition probabilities $\{P(\cdot|s,a)\}$ and the agent’s estimates $\{P_{e}(\cdot|s,a)\}$.
Lemma 2.
Execution of Q-UCRL (Algorithm 2) up to the beginning of epoch $e+1$ with total $e$ completed epochs comprising $t=1,2,\ldots,t_{e}$ rounds, guarantees that the following holds for $\tilde{P}_{e+1}$:
$\mathbb{P}[|\tilde{P}_{e+1}(s^{\prime}|s,a)-P(s^{\prime}|s,a)|\leq{\frac{C\log(S/\delta)}{\nu_{e}(s,a)}}]\geq 1-\delta$ (27)
$\forall(s,a,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}$, where $C={c\log^{1/2}(T\sqrt{S})}$ for some $c\in\mathbb{R}$ defined in Eq. (25) and $\{\nu_{e}(s,a)\}$ are the state-action pair visitation counts up in epoch $e$, as depicted in Line 12 of Algorithm 2, and $\{P(s^{\prime}|s,a),\tilde{P}_{e+1}(s^{\prime}|s,a)\}$ are the actual and estimated transition probabilities by Eq. (21).
Proof.
For all $(s,a)\in\mathcal{S}\times\mathcal{A}$ , since $\tilde{P}_{e+1}$ is obtained by Eq. (21), we obtain using Lemma 1:
$\mathbb{P}[{||\tilde{P}_{e+1}(\cdot|s,a)-P(\cdot|s,a)||_{\infty}\leq\frac{\log(S/\delta)}{n_{e+1}^{s_{i},a_{i}}}}]\geq 1-\delta,$ (28)
Thus, by substituting $n_{e}^{s_{i},a_{i}}$ with $\nu_{e}{(s,a)}/C$, Eq. (28) would produce:
${\mathbb{P}[||\tilde{P}_{e+1}(\cdot|s,a)-P(\cdot|s,a)||_{\infty}\leq\frac{C\log(S/\delta)}{\nu_{e}(s,a)}]\geq 1-\delta}$ (29)
Eq. (29) can be transformed into Eq. (27) by taking the entry-wise expression.
Lemma 3.
Execution of Q-UCRL (Algorithm 2) up to the beginning of epoch $e+1$ with total $e$ completed epochs comprising $t=1,2,\ldots,t_{e}$ rounds, guarantees that the following holds for $\hat{P}_{e+1}$:
$\mathbb{P}[|\hat{P}_{e+1}(s^{\prime}|s,a)-P(s^{\prime}|s,a)|\leq\frac{eC\log(e{S}/\delta)}{N_{e+1}(s,a)}]\geq 1-\delta$ (30)
$\forall(s,a,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}$, where $C={c\log^{1/2}(T\sqrt{S})}$ for some $c\in\mathbb{R}$ and $\{\nu_{e}(s,a)\}$ are the state-action pair visitation counts up in epoch $e$, as depicted in Line 12 of Algorithm 2, and $\{P(s^{\prime}|s,a),\hat{P}_{e+1}(s^{\prime}|s,a)\}$ are the actual and weighted average of transition probabilities by Eq. (22).
Proof.
For all $(s,a,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}$ , $\hat{P}_{e+1}(s^{\prime}|s,a)$ could be equivalently expressed as:
$\hat{P}_{e+1}(s^{\prime}|s,a)=\frac{\sum_{i=1}^{e}\nu_{i}(s,a)\tilde{P}_{i}(s^{\prime}|s,a)}{\sum_{j=1}^{e}\nu_{j}(s,a)}$ (31)
Denote event $|\tilde{P}_{i}(s^{\prime}|s,a)-P(s^{\prime}|s,a)|\leq{\frac{C\log(S/\delta)}{\nu_{e}(s,a)}}$ as event $E_{i}$. Thus if events $E_{1},...,E_{e}$ all occur, we would have:
$|\hat{P}_{e+1}(s^{\prime}|s,a)-P(s^{\prime}|s,a)|$ (32)
$=\Bigg{|}\frac{\sum_{i=1}^{e}\nu_{i}(s,a)(\tilde{P}_{i}(s^{\prime}|s,a)-P(s^{\prime}|s,a))}{\sum_{j=1}^{e}\nu_{j}(s,a)}\Bigg{|},$ (33)
$\leq\frac{\sum_{i=1}^{e}\nu_{i}(s,a)|\tilde{P}_{i}(s^{\prime}|s,a)-P(s^{\prime}|s,a)|}{\sum_{j=1}^{e}\nu_{j}(s,a)},$ (34)
$\leq\frac{\sum_{i=1}^{e}C\log({S}/\delta)}{\sum_{j=1}^{e}\nu_{j}(s,a)},$ (35)
$=\frac{eC\log({S}/\delta)}{\sum_{j=1}^{e}\nu_{j}(s,a)},$ (36)
$=\frac{eC\log({S}/\delta)}{N_{e+1}(s,a)},$ (37)
Where Eq. (34) is the result of triangle inequality and Eq. (35) is by Lemma 2. The probability of events $E_{1},...,E_{e}$ all occurring is:
$\mathbb{P}[E_{1}\cap...\cap E_{e}],$ (38)
$=1-\mathbb{P}[E_{1}^{C}\cup...\cup E_{e}^{C}],$ (39)
$\geq 1-\sum_{i=1}^{e}\mathbb{P}[E_{i}^{C}],$ (40)
$\geq 1-e\delta$ (41)
Where Eq. (41) is by Lemma 2. By combing Eq. (37) and Eq. (41) and substituting $e\delta$ with $\delta$, we would obtain Eq. (30) ∎
Lemma 4.
Execution of Q-UCRL (Algorithm 2) up to total $e$ epochs comprising $t=1,2,\ldots,t_{e}$ rounds, guarantees that the following holds for confidence parameter $\delta_{e}=\frac{1}{SAt_{e}^{7}}$ with probability at least $1-1/t_{e}^{6}$:
$\|{P}_{e}(\cdot|s,a)-{P}(\cdot|s,a)\|_{1},$
$\leq\frac{7S(1+Ce)\log(S^{2}AT)+SCelog{(eS)}}{\max\{1,N_{e}(s,a)\}},$ (42)
$\forall(s,a)\in\mathcal{S}\times\mathcal{A}$, where $\{N_{e-1}^{(s,a)}\}$ is the cumulative state-action pair visitation counts up to epoch $e$, as depicted in Line 17 of Algorithm 2, and $\{P(\cdot|s,a),P_{e}(\cdot|s,a)\}$ are the actual and estimated transition probabilities.
Proof.
To prove the claim in Eq. (42), note that for an arbitrary pair $(s,a)$ we have:
$\|{P}_{e}(\cdot|s,a)-{P}(\cdot|s,a)\|_{1}$ (43)
$\leq$ $\|{P}_{e}(\cdot|s,a)-\hat{P}_{e}(\cdot|s,a)\|_{1}$
$+\|\hat{P}_{e}(\cdot|s,a)-{P}(\cdot|s,a)\|_{1},$
$\leq$ $\frac{7S\log(S^{2}At_{e})}{\max\{1,N_{e}(s,a)\}}+\|\hat{P}_{e}(\cdot|s,a)-{P}(\cdot|s,a)\|_{1},$ (44)
$=$ $\frac{7S\log(S^{2}At_{e})}{\max\{1,N_{e}(s,a)\}}$ (45)
$+\sum_{s^{\prime}\in\mathcal{S}}|\underbrace{\hat{P}_{e}(s^{\prime}|s,a)-{P}(s^{\prime}|s,a)}_{\text{(a)}}|,$
where Eq. (44) is a direct consequence of constraint Eq. (19) of optimization problem Eq. (15). Next, in order to analyze (a) in Eq. (45), consider the following event:
$\mathcal{E}$ =$=\Bigg{\{}|{\hat{P}_{e}(s^{\prime}|s,a)-{P}(s^{\prime}|s,a)}|$
$\leq\frac{7\log(S^{2}At_{e})}{\max\{1,N_{e}(s,a)\}},\forall(s,a,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}\Bigg{\}},$ (46)
For a tuple $(s,a,s^{\prime})$, by replacing $\delta=e^{-\epsilon\cdot N_{e}(s,a)}$ into Eq. (1) from Lemma 1, we get:
$\mathbb{P}\left[|{\hat{P}_{e}(s^{\prime}|s,a)-{P}(s^{\prime}|s,a)}|\leq\frac{Ce\log(e{S})}{N_{e}{(s,a)}}+Ce\epsilon\right]$
$\quad\quad\geq 1-e^{-\epsilon\cdot N_{e}(s,a)},$ (47)
By plugging in $\epsilon=\frac{\log(S^{2}At_{e}^{7})}{N_{e}(s,a)}$ into Eq. (47), we get:
$\mathbb{P}\left[|{\hat{P}_{e}(s^{\prime}|s,a)-{P}(s^{\prime}|s,a)}|\geq\frac{Ce\log{(eS)}+Ce\log(S^{2}At_{e}^{7})}{N_{e}(s,a)}\right]$
$\quad\quad\leq\frac{1}{S^{2}At_{e}^{7}},$ (48)
Finally, we sum over all possible values of $N_{e}(s,a)$ as well as tuples $(s,a,s^{\prime})$ to bound the probability in the RHS of (48) that implies failure of event $\mathcal{E}$:
$\sum_{(s,a,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}}\sum_{N_{e}(s,a)=1}^{t_{e}}\frac{1}{S^{2}At_{e}^{7}}\leq\frac{1}{t_{e}^{6}}$ (49)
Here, it is critical to highlight that we obtain this high probability bound for model errors by careful utilization of Lemma 1 and quantum signals $\bm{\psi}^{t}$ via Eq. (42) - (49), while avoiding classical concentration bounds for probability norms that use martingale style analysis such as in the work (?). Finally, plugging the bound of term (a) as described by event $\mathcal{E}$, we obtain the following with probability $1-1/t_{e}^{6}$:
$\|{P}_{e}(\cdot|s,a)-{P}(\cdot|s,a)\|_{1}$
$\leq\frac{7S\log(S^{2}At_{e})}{\max\{1,N_{e}(s,a)\}}+\sum_{s^{\prime}\in\mathcal{S}}\frac{Ce\log{(eS)}+Ce\log(S^{2}At_{e}^{7})}{\max\{1,N_{e}(s,a)\}},$ (50)
$\leq\frac{7S\log(S^{2}At_{e})}{\max\{1,N_{e}(s,a)\}}+\sum_{s^{\prime}\in\mathcal{S}}\frac{Ce\log{(eS)}+7Ce\log(S^{2}At_{e})}{\max\{1,N_{e}(s,a)\}},$ (51)
$=\frac{7S(1+Ce)\log(S^{2}At_{e})+SCelog{(eS)}}{\max\{1,N_{e}(s,a)\}},$ (52)
$\leq\frac{7S(1+Ce)\log(S^{2}AT)+SCelog{(eS)}}{\max\{1,N_{e}(s,a)\}}.$ (53)
This proves the claim as in the statement of the Lemma. ∎
Remark 2.
Note that for probability vectors ${P}_{e}(\cdot|s,a),{P}(\cdot|s,a)$ we have the bound: $\|{P}_{e}(\cdot|s,a)-{P}(\cdot|s,a)\|_{1}\leq 2S$, in the worst case scenario. However, the probability that a bad event (i.e., complementary of Eq. (42)) happens is at most $\frac{1}{t_{e}^{6}}$.
Regret Decomposition: Regret of Q-UCRL defined in Eq. (10) can be decomposed as follows:
$\mathcal{R}_{[0,T-1]}$ $=T.\Gamma_{\pi^{*}}^{P}-\mathbb{E}\Big{[}\sum_{t=0}^{T-1}r(s_{t},a_{t})\Big{]},$ (54)
$=T.\Gamma_{\pi^{*}}^{P}-\sum_{e=1}^{E}T_{e}\Gamma_{\pi^{*}_{e}}^{P}$
$+\sum_{e=1}^{E}T_{e}\Gamma_{\pi^{*}_{e}}^{P}-\mathbb{E}\Big{[}\sum_{t=0}^{T-1}r(s_{t},a_{t})\Big{]},$ (55)
$=\sum_{e=1}^{E}T_{e}\big{(}\underbrace{\Gamma_{\pi^{*}}^{P}-\Gamma_{\pi^{*}_{e}}^{P}}_{\text{(a)}}\big{)}$
$+\sum_{e=1}^{E}T_{e}\Gamma_{\pi^{*}_{e}}^{P}-\mathbb{E}\Big{[}\sum_{t=0}^{T-1}r(s_{t},a_{t})\Big{]},$ (56)
$=\sum_{e=1}^{E}T_{e}\Gamma_{\pi^{*}_{e}}^{P}-\mathbb{E}\Big{[}\sum_{t=0}^{T-1}r(s_{t},a_{t})\Big{]},$ (57)
$\leq\sum_{e=1}^{E}T_{e}\Gamma_{\pi_{e}}^{{P}_{e}}-\mathbb{E}\Big{[}\sum_{t=0}^{T-1}r(s_{t},a_{t})\Big{]},$ (58)
$=\sum_{e=1}^{E}T_{e}\Gamma_{\pi_{e}}^{{P}_{e}}-\sum_{e=1}^{E}T_{e}\Gamma_{\pi_{e}}^{{P}}$
$+\sum_{e=1}^{E}T_{e}\Gamma_{\pi_{e}}^{{P}}-\mathbb{E}\Big{[}\sum_{t=0}^{T-1}r(s_{t},a_{t})\Big{]},$ (59)
$=\sum_{e=1}^{E}T_{e}\Gamma_{\pi_{e}}^{{P}_{e}}-\sum_{e=1}^{E}T_{e}\Gamma_{\pi_{e}}^{{P}}$
$+\Big{[}\sum_{e=1}^{E}\sum_{t=t_{e-1}+1}^{t_{e}}\underbrace{\Gamma_{\pi_{e}}^{{P}}-\mathbb{E}[r(s_{t},a_{t})]}_{\text{(}b)}\Big{]},$ (60)
$=\sum_{e=1}^{E}T_{e}\Big{(}\Gamma_{\pi_{e}}^{\tilde{P}_{e}}-\Gamma_{\pi_{e}}^{{P}}\Big{)}.$ (61)
Firstly, (a) in Eq. (56) is 0 owing to the fact that the agent would have gotten $\pi_{e}^{*}$ exactly as the true $\pi^{*}$ in epoch $e$ solving for original problem Eq. (11), had it known the true transition model $P$. Next, Eq. (58) is because $\{P_{e},\pi_{e}\}$ are outputs from the optimistic problem Eq. (15) solved at round $e$, which extends the search space of Eq. (11) and therefore always upper bounds true long-term rewards. Finally, term (b) in Eq. (60) is 0 in expectation, because the RL agent actually collects rewards by playing policy $\pi_{e}$ against the true $P$ during epoch $e$.
Regret in terms of Bellman Errors: Here, we introduce the notion of Bellman errors $B_{\pi}^{{P}_{e}}(s,a)$ at epoch $e$ as follows:
$B_{\pi}^{{P}_{e}}(s,a)\triangleq\lim_{\gamma\rightarrow 1}$ $~{}[Q_{\pi}^{{P}_{e}}(s,a;\gamma)-r(s,a)$
$-\gamma\sum_{s^{\prime}\in\mathcal{S}}P(s^{\prime}|s,a)V_{\pi}^{{P}_{e}}(s;\gamma)],$ (62)
which essentially captures the gap in the average rewards by playing action $a$ in state $s$ accordingly to some policy $\pi$ in one step of the optimistic transition model $P_{e}$ w.r.t. the actual probabilities $P$. Also, we note here that the definition of $Q_{\pi}^{{P}_{e}}$ function is associated to the value function definition in eq (5).
This leads us to the following key results that connect the regret expression in Eq. (61) and the Bellman errors.
Lemma 5.
[ Lemma 5.2, (?)] The difference in long-term expected reward by playing optimistic policy of epoch $e$, i.e., $\pi_{e}$ on the optimistic transition model estimates, i.e., $\Gamma_{\pi_{e}}^{{P}_{e}}$, and the expected reward collected by playing $\pi_{e}$ on the actual model $P$, i.e., $\Gamma_{\pi_{e}}^{{P}}$ is the long-term average of the Bellman errors. Mathematically, we have:
$\Gamma_{\pi_{e}}^{{P}_{e}}-\Gamma_{\pi_{e}}^{{P}}=\sum_{s,a}\rho_{\pi_{e}}^{P}(s,a)B_{\pi_{e}}^{{P}_{e}}(s,a),$ (63)
where $B_{\pi}^{{P}_{e}}(s,a)$ is defined in Eq. (62).
With Eq. (62), it necessary to mathematically bound the Bellman errors. To this end, we leverage the following result that characterizes the Bellman errors in terms of $\tilde{h}(\cdot)$, a quantity that measures the bias contained in the optimistic MDP model.
Lemma 6.
[Lemma 5.3, (?)] The Bellman errors for any arbitrary state-action pair $(s,a)$ corresponding to epoch $e$ is bounded as follows:
$B^{\pi_{e},{P}_{e}}(s,a)\leq\Big{\|}{P}_{e}(\cdot|s,a)-{P}(\cdot|s,a)\Big{\|}_{1}\|\tilde{h}(\cdot)\|_{\infty},$ (64)
where $\tilde{h}(\cdot)$ (defined in (?)) is the bias of the optimistic MDP $P_{e}$ for which the optimistic policy $\pi_{e}$ obtains the maximum expected reward in the confidence interval.
Consequently, by plugging in the bound of the error arising from estimation of transition model in Lemma 4 into eq 64, we obtain the following result with probability $1-1/t_{e}^{6}$:
$B^{\pi_{e},{P}_{e}}(s,a)$
$\leq\frac{7S(1+Ce)\log(S^{2}AT)+SCelog{(eS)}}{\max\{1,N_{e}(s,a)\}}\|\tilde{h}(\cdot)\|_{\infty},$ (65)
Consequently, it is necessary to bound the model bias error term in Eq. (65). We recall that the optimistic MDP $P_{e}$ corresponding to epoch $e$ maximizes the average reward in the confidence set as described in the formulation of eq (15). In other words, $\Gamma_{\pi_{e}}^{P_{e}}\geq\Gamma_{\pi_{e}}^{P^{\prime}}$ for every $P^{\prime}$ in the confidence interval. Specifically, the model bias for optimistic MDP can be explicitly defined in terms of Bellman equation (?), as follows:
$\tilde{h}(s)$ $\triangleq r_{\pi_{e}}(s)-\lambda_{\pi_{e}}^{P_{e}}+\Big{(}P_{e,\pi_{e}}(\cdot|s)\Big{)}^{T}\tilde{h}(\cdot),$ (66)
where the following simplified notations are used: $r_{\pi_{e}}(s)=\sum_{a\in\mathcal{A}}\pi_{e}(a|s)r(s,a)$, $P_{e,\pi_{e}}(s^{\prime}|s)=\sum_{a\in\mathcal{A}}\pi_{e}(a|s)P_{e}(s^{\prime}|s,a)$. Next, we incorporate the following result that correlates model bias $\tilde{h}(\cdot)$ to the mixing time $T_{\texttt{mix}}$ of MDP $\mathcal{M}$.
Lemma 7.
[Lemma D.5, (?)] For an MDP with the transition model $P_{e}$ that generates rewards $r(s,a)$ on playing policy $\pi_{e}$, the difference of biases for states $s$ and $s^{\prime}$ are bounded as:
$\tilde{h}(s)-\tilde{h}(s^{\prime})\leq T_{\texttt{mix}},\forall s,s^{\prime}\in\mathcal{S}.$ (67)
Plugging in the bound of bias term $\tilde{h}(\cdot)$ into Eq. (65), we get the final bound for Bellman error with probability $1-1/T^{6}$ as:
$B^{\pi_{e},{P}_{e}}(s,a)$
$\leq\frac{7S(1+Ce)\log(S^{2}AT)+SCelog{(eS)}}{\max\{1,N_{e}(s,a)\}}T_{\texttt{mix}}.$ (68)
Regret Bound for Q-UCRL We now present the final bound for regret incurred by Q-UCRL over a $T$ length training horizon.
Theorem 1.
In an unknown MDP environment $\mathcal{M}\triangleq(\mathcal{S},\mathcal{A},P,r,D)$, the regret incurred by Q-URL (Algorithm 2) across $T$ rounds is bounded as follows:
$\mathcal{R}_{[0,T-1]}=\mathcal{O}\Bigg{(}{S^{5}A^{4}T_{\texttt{mix}}{\log^{3}(\frac{T}{SA})\log^{1/2}(T\sqrt{S})\log(S^{2}AT)}}\Bigg{)}.$ (69)
Proof.
Using the final expression from regret decomposition, i.e., Eq. (61), we have:
$\mathcal{R}_{[0,T-1]}$
$=\sum_{e=1}^{E}T_{e}\Big{(}\Gamma_{\pi_{e}}^{\tilde{P}_{e}}-\Gamma_{\pi_{e}}^{{P}}\Big{)},$ (70)
$=\frac{1}{T}\sum_{e=1}^{E}\sum_{t=t_{e-1}+1}^{t_{e}}\sum_{s,a}\rho_{\pi_{e}}^{P}B^{\pi_{e},{P}_{e}}(s,a),$ (71)
$\leq\sum_{e=1}^{E}\sum_{t=t_{e-1}+1}^{t_{e}}\frac{7S(1+Ce)\log(S^{2}AT)+SCelog{(eS)}}{\max\{1,N_{e}(s_{t},a_{t})\}}T_{\texttt{mix}},$ (72)
$=\sum_{e=1}^{E}\sum_{t=t_{e-1}+1}^{t_{e}}\sum_{s,a}\Big{(}\mathbf{1}[s_{t}=s,a_{t}=a]\cdot$
$\hskip 28.45274pt\frac{7S(1+Ce)\log(S^{2}AT)+SCelog{(eS)}}{\max\{1,N_{e}(s,a)\}}T_{\texttt{mix}}\Big{)},$ (73)
$=\sum_{e=1}^{E}\sum_{s,a}\nu_{e}(s,a)\cdot$
$\hskip 28.45274pt\frac{7S(1+Ce)\log(S^{2}AT)+SCelog{(eS)}}{\max\{1,N_{e}(s,a)\}}T_{\texttt{mix}},$ (74)
$=\sum_{s,a}T_{\texttt{mix}}(7S(E+CE^{2})\log(S^{2}AT)+SCE^{2}\log{(ES)})$
$\cdot\sum_{e=1}^{E}\frac{\nu_{e}(s,a)}{\max\{1,N_{e}(s,a)\}},$ (75)
$\leq\sum_{s,a}T_{\texttt{mix}}(7S(E+CE^{2})\log(S^{2}AT)+SCE^{2}\log{(ES)})E,$
$=T_{\texttt{mix}}E(7S^{2}A(E+CE^{2})\log(S^{2}AT)+S^{2}ACE^{2}\log{(ES)}),$
$=\mathcal{O}\Bigg{(}{S^{5}A^{4}T_{\texttt{mix}}{\log^{3}(\frac{T}{SA})\log^{1/2}(T\sqrt{S})\log(S^{2}AT)}}\Bigg{)}.$ (76)
where in Eq. (71), we directly incorporate the result of Lemma 5. Next, we obtain Eq. (72) by using the Bellman error bound in Eq. (68), as well as unit sum property of occupancy measures (?). Furthermore, in Eq. (72) note that we have omitted the regret contribution of bad event (Remark 2) which is atmost $\tilde{\mathcal{O}}\Bigg{(}\frac{2ST_{\texttt{mix}}}{T^{4}}\Bigg{)}$ as obtained below:
$\sum_{e=1}^{E}$ $\sum_{t=t_{e-1}+1}^{t_{e}}\frac{2ST_{\texttt{mix}}}{t_{e}^{6}}\leq\sum_{e=1}^{E}\frac{2ST_{\texttt{mix}}}{t_{e}^{5}},$ (77)
$\leq\sum_{t_{e}=1}^{T}\frac{2ST_{\texttt{mix}}}{t_{e}^{5}}=\tilde{\mathcal{O}}\Bigg{(}\frac{2ST_{\texttt{mix}}}{T^{4}}\Bigg{)}.$ (78)
Subsequently, Eq. (74) - (75) are obtaining by using the definition of epoch-wise state-action pair visitation frequencies $\nu_{e}(s,a)$ and the epoch termination trigger condition in Line 13 of Algorithm 2. Finally, we obtain Eq. (76) by the using Proposition 18 in Appendix C.2 of (?). ∎
Remark 3.
We reiterate that our characterization of regret incurred by Q-UCRL in Theorem 1 is a martingale-free analysis and thus deviates from its classical counterparts (?, ?, ?).
6 Conclusions and Future Work
This paper demonstrates that quantum computing helps provide significant reduction in the regret bounds for infinite horizon reinforcement learning with average rewards. This paper not only unveils the quantum potential for bolstering reinforcement learning but also beckons further exploration into novel avenues that bridge quantum and classical paradigms, thereby advancing the field to new horizons.
We note that this paper considers a model-based setup, while model-free setup for the problem is an important problem for the future.
A Convexity of Optimization Problem
In order to solve the optimistic optimization problem in Eq. (15) - (19) for epoch $e$, we adopt the approach proposed in (?, ?). The key is that the constraint in (18) seems non-convex. However, that can be made convex with an inequality as
$\sum_{a\in\mathcal{A}}\rho(s^{\prime},a)\leq\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}{P}_{e}(s^{\prime}|s,a)\rho(s,a),~{}\forall s^{\prime}\in\mathcal{S}$ (79)
This makes the problem convex since $xy\geq c$ is convex region in the first quadrant.
However, since the constraint region is made bigger, we need to show that the optimal solution satisfies (18) with equality. This follows since
$\sum_{s^{\prime}\in\mathcal{S}}\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}{P}_{e}(s^{\prime}|s,a)\rho(s,a)=1=\sum_{s^{\prime}\in\mathcal{S}}\sum_{a\in\mathcal{A}}\rho(s^{\prime},a)$ (80)
Thus, since we have point-wise inequality in (79) with the sum of the two sides being the same, thus, the equality holds for all $s^{\prime}$.
B A brief overview of Quantum Amplitude Amplification
Quantum amplitude amplification stands as a critical outcome within the realm of quantum mean estimation. This pivotal concept empowers the augmentation of the amplitude associated with a targeted state, all the while suppressing the amplitudes linked to less desirable states. The pivotal operator, denoted as $Q$, embodies the essence of amplitude amplification: $Q=2|\psi\rangle\langle\psi|-I$, where $|\psi\rangle$ signifies the desired state and $I$ represents the identity matrix. This operator orchestrates a double reflection of the state $|\psi\rangle$ initially about the origin and subsequently around the hyperplane perpendicular to $|\psi\rangle$ culminating in a significant augmentation of the amplitude of $|\psi\rangle$.
Moreover, the application of this operator can be iterated to achieve a repeated amplification of the desired state’s amplitude, effectively minimizing the amplitudes of undesired states. Upon applying the amplitude amplification operator a total of $t$ times, the outcome is $Q^{t}$, which duly enhances the amplitude of the desired state by a scaling factor of $\sqrt{N}/M$, where $N$ corresponds to the overall count of states and $M$ signifies the number of solutions.
The implications of quantum amplitude amplification extend across a diverse spectrum of applications within quantum algorithms. By bolstering their efficacy and hastening the resolution of intricate problems, quantum amplitude amplification carves a substantial niche. Noteworthy applications encompass:
  1. 1.
    Quantum Search: In the quantum algorithm for unstructured search, known as Grover’s algorithm (?), the technique of amplitude amplification is employed to enhance the amplitude of the target state and decrease the amplitude of the non-target states. As a result, this technique provides a quadratic improvement over classical search algorithms.
  2. 2.
    Quantum Optimization: Quantum amplification can be used in quantum optimization algorithms, such as Quantum Approximate Optimization Algorithm (QAOA) (?), to amplify the amplitude of the optimal solution and reduce the amplitude of sub-optimal solutions. This can lead to an exponential speedup in solving combinatorial optimization problems.
  3. 3.
    Quantum Simulation: Quantum amplification has also been used in quantum simulation algorithms, such as Quantum Phase Estimation (QPE) (?), to amplify the amplitude of the eigenstate of interest and reduce the amplitude of other states. This can lead to an efficient simulation of quantum systems, which is an intractable problem in the classical domain.
  4. 4.
    Quantum Machine Learning: The utilization of quantum amplification is also evident in quantum machine learning algorithms, including Quantum Support Vector Machines (QSVM) (?). The primary objective is to amplify the amplitude of support vectors while decreasing the amplitude of non-support vectors, which may result in an exponential improvement in resolving particular classification problems.
In the domain of quantum Monte Carlo methods, implementing quantum amplitude amplification can enhance the algorithm’s efficiency by decreasing the amount of samples needed to compute the integral or solve the optimization problem. By amplifying the amplitude of the desired state, a substantial reduction in the number of required samples can be achieved while still maintaining a certain level of accuracy, leading to significant improvements in efficiency when compared to classical methods. This technique has been particularly useful in (?) for achieving efficient convergence rates in quantum Monte Carlo simulations.